OpenStudy (anonymous):

The sum of the digits of a certain two-digit number is 7. Reversing its digits increases the number by 9. What is the number?

hero (hero):

34 Luis, what are you smokin?

hero (hero):

@Luis_Rivera

hero (hero):

You're interpreting it wrong

hero (hero):

Why would it say "increases the number by 9" rather than "7 increases by 9"?

hero (hero):

It is referring to a number that we don't yet know

hero (hero):

No, you're interpreting it wrong, you just don't realize it

hero (hero):

I'm sorry that you don't understand that your steps are flawed.

hero (hero):

only one person deserves medals

hero (hero):

34 43 3 + 4 = 7 34 + 9 = 43 (reversed digits)

hero (hero):

@myininaya

hero (hero):

@satellite73

hero (hero):

@Zarkon

OpenStudy (anonymous):

i don't think 61 is 9 more than 16, so it is unlikely that 16 and 61 is the answer, at least not the way i read the question

myininaya (myininaya):

Hey Hero... I know you mean well, but it comes off a little mean to tell someone they don't deserve medals.

hero (hero):

myininaya , you always assume I'm being mean. Can I get the benefit of the doubt at least once?

myininaya (myininaya):

So anyways how I interpret the problem is... The sum of the digits of a certain two-digit number is 7. Reversing its digits increases the number by 9. What is the number? Let d1d2 be a two-digit number (d1d2=d1*10+d2*1) d1+d2=7 d2d1=9+d1d2 =>10*d2+d1=9+d1*10+d2 Is how I read the equation I could be wrong... The last sentence doesn't make sense because it is talking about two different numbers.

myininaya (myininaya):

The last sentence acts as if it is talking about the same numbers. Changing the order of the digits clearly gives us a difference number unless d1=d2 which it cannot since d1+d2=7 And there is no integer a such that a+a=7

myininaya (myininaya):

9d2-9d1=9 => d2-d1=1 + (d2+d1=7) ---------- 2d2=8 d2=4 => d2-d1=1=> 4-d1=1 =>d1=3 So the digit number that I come up with is d1d2=34

hero (hero):

very nice

myininaya (myininaya):

Thanks Hero.

OpenStudy (vishal_kothari):

a + b = 7 10a + b + 9 = 10b + a 9a - 9b = -9 9a = 9b - 9 a = b - 1 Plugging back into the original equation: (b - 1) + b = 7 2b - 1 = 7 2b = 8 b = 4 a = 3 34.

OpenStudy (anonymous):

\(34\) it is, \(3+4=7\) and \(43-34=9 \).